OpenStudy (anonymous):

what is the derivative of square root of x-5

OpenStudy (mayankdevnani):

what do you think ??

OpenStudy (anonymous):

x^1/2 -5^1/2?

OpenStudy (mayankdevnani):

no,the answer would be :- \[\large \bf \frac{1}{2 \times \sqrt{x-5} }\]

OpenStudy (anonymous):

ohhhh ok .. i got thanks

OpenStudy (mayankdevnani):

Steps to solve :-

OpenStudy (mayankdevnani):

hope you understand. @forzamilan

OpenStudy (mayankdevnani):

STEP BY STEP SOLUTION :-